Florida Groves has 380 bonds outstanding that are selling at 97.4 percent of face value. The

firm also has 2,600 shares of preferred stock valued at $61 a share and 37,500 shares of common

stock valued at $19 a share. What weight should be assigned to the common stock when

computing the weighted average cost of capital?

=57.40%

-Please check the answers and show all work typed out. No excel or grid style please as I am on mobile.
Answer #1

Total Value of bond =380*1000*97.4%=$370120

Total Value of Equity=37500*19=$712500

Total Value of preferred stock=2600*61=$158600

Total Value =$370120+712500+158600=$1241220

weight assigned to the common stock=Total Value of Equity/Total Value

=$712500/1241220=57.40%